# Robustness in Large Language Models: A Survey of Mitigation Strategies and Evaluation Metrics

### v1

@@ 7 Challenges and Future work
    Causal and Invariant Learning: Moving beyond correlational learning towards models that understand underlying causal relationships or learn representations that are inherently invariant to changes in domain or distribution (Ye et al., 2024). This could lead to more fundamentally robust models less reliant on spurious cues.
@@
    Hybrid Approach: Exploring hybrid approaches that combine multiple processing steps, as well as investigating methods to optimize their computational efficiency, are promising directions for advancing the field and ensuring the safe and reliable deployment of large language models in a wide range of real-world applications.
@@
    Efficient Robust Training and Alignment: Developing more scalable adversarial training methods (e.g., improving continuous AT), more efficient robust optimization techniques, and alignment methods (like RLHF variants) that explicitly minimize the alignment tax or are inherently more robust to reward model imperfections (e.g., reward-robust RLHF) (Yan et al., 2024b).
@@
    Robustness in Multimodal and Agentic Systems: Extending robustness research to the unique challenges of MLLMs and emerging LLM-based agent systems, considering factors like tool use, memory, and interaction with environments (Jiang et al., 2025; Yu et al., 2025).
@@
    Interpretability for Robustness: Leveraging interpretability techniques to better understand why models fail in specific robustness scenarios (e.g., identifying responsible components (Yang et al., 2025b) or internal representations of truth/lies (Bürger et al., 2024)) and using these insights to design targeted interventions or more robust architectures like a reconstruction attack (Stacey et al., 2024; 2022).
In conclusion, while LLMs have demonstrated transformative potential, ensuring their robustness is paramount for realizing this potential responsibly. The field has made substantial progress in identifying vulnerabilities and developing initial solutions, but significant challenges remain. Future work focusing on fundamental understanding (causality, invariance), efficient and scalable robust training methods, comprehensive and adaptive evaluation, and extending robustness considerations to newer paradigms like MLLMs and agents will be crucial for building truly reliable and trustworthy AI systems required for widespread societal integration.
While LLMs have demonstrated transformative capabilities, ensuring their robustness remains critical for safe and responsible real-world adoption. Though the field has advanced significantly in diagnosing vulnerabilities and proposing mitigation strategies, fundamental challenges persist. Future progress will depend on four key priorities: (1) deepening theoretical insights into causality and invariance, (2) developing efficient robust training frameworks that scale with model size, (3) creating adaptive evaluation systems that address evolving failure modes, and (4) extending robustness principles to emerging paradigms like multimodal LLMs (MLLMs) and autonomous AI agents. Addressing these challenges systematically will be essential to transition from promising prototypes to trustworthy, reliably deployable AI systems.

### v2

@@ 6.2 Future Research Directions
    Causal and Invariant Learning: Moving beyond correlational learning towards models that understand underlying causal relationships or learn representations that are inherently invariant to changes in domain or distribution (ye2024spuriouscorrelationsmachinelearning). This could lead to more fundamentally robust models less reliant on spurious cues.
@@
    Hybrid Approach: Exploring hybrid approaches that combine multiple processing steps, as well as investigating methods to optimize their computational efficiency, are promising directions for advancing the field and ensuring the safe and reliable deployment of LLMs in a wide range of real-world applications.
@@
    Efficient Robust Training and Alignment: Developing more scalable adversarial training methods (e.g., improving continuous AT), more efficient robust optimization techniques, and alignment methods (like RLHF variants) that explicitly minimize the alignment tax or are inherently more robust to reward model imperfections (e.g., reward-robust RLHF) (yan2024rewardrobustrlhfllms).
@@
    Robustness in Multimodal and Agentic Systems: Extending robustness research to the unique challenges of MLLMs and emerging LLM-based agent systems, considering factors like tool use, memory, and interaction with environments (jiang2025surveyadversarialrobustnessmultimodal; yu2025surveytrustworthyllmagents).
@@
    Interpretability for Robustness: Leveraging interpretability techniques to better understand why models fail in specific robustness scenarios (e.g., identifying responsible components (yang2025enhancingsemanticconsistencylarge) or internal representations of truth/lies (burger2024truth)) and using these insights to design targeted interventions or more robust architectures like a reconstruction attack (stacey2024atomicinferencenligenerated; stacey2022logicalreasoningspanlevelpredictions).
In conclusion, while LLMs have demonstrated transformative capabilities, ensuring their robustness remains critical for safe and responsible real-world adoption. Though the field has advanced significantly in diagnosing vulnerabilities and proposing mitigation strategies, fundamental challenges persist. Future progress will depend on four key priorities: (1) deepening theoretical insights into causality and invariance, (2) developing efficient and robust training frameworks that scale with model size, (3) creating adaptive evaluation systems that address evolving failure modes, and (4) extending robustness principles to emerging paradigms like MLLMs and autonomous AI agents. Addressing these challenges systematically will be essential to transition from promising prototypes to trustworthy, reliably deployable AI systems.